1: A hydrocarbon contains 7.7% by mass of carbon. The relative mass of the compound is 78. Derive the empirical formula of the compound, it's molecular formula, name the hydrogcarbons and write the structural formula [H=1,C=12]


2: A is a dilute tetraoxpsulphate VI acid, B contains 8g/dm³ of sodium hydroxide of solution. 25cm³ of B requires 20cm³ of A for complete neutralization. Calculate the concentration of B and A in mol/dm³, and, the number of hydrogen ions in 1.0dm³od solution A.


1
Expert's answer
2021-09-20T02:24:49-0400

7.7/100 × 78 = 6

78-6 = 72/12 = 6

Therefore the structural formula is

C6H6